OpenStudy (anonymous):

Can someone help me solve ln(x)-ln(x+1)=ln(x+3)-ln(x+5)

OpenStudy (anonymous):

are you trying to solve for x?

OpenStudy (anonymous):

I believe so. The instructions just say solve giving exact solution

OpenStudy (jennychan12):

properties of natural logs. ln(x(x+1)) = lnI(x+5)(x+3) go from there.

OpenStudy (anonymous):

- means division

OpenStudy (anonymous):

\[\ln(\frac{ x }{ x+1 }) = \ln(\frac{ x+3 }{ x+5 })\]

OpenStudy (jennychan12):

oh whoops sorry. i haven't done alg 2 in years!

OpenStudy (anonymous):

yeah see that's what I thought it was too Veritas, but then there I got stuck

OpenStudy (anonymous):

no problem. Yeah I'm trying to see how this would even help to solve the equation though.

OpenStudy (jennychan12):

take e of both sides to cancel out the ln's

OpenStudy (anonymous):

no this is the natural log

OpenStudy (anonymous):

Ignore that lat bit from me. :)

OpenStudy (anonymous):

yeah I think Jenny's answer is the way to solve.

OpenStudy (anonymous):

ok I'm gonna try and solve now

OpenStudy (anonymous):

\[e(\ln (\frac{ x }{ x+1 }))=e(\ln(\frac{ x+3 }{ x+5 }))\]

OpenStudy (anonymous):

and those e's should be raised to the power of the natural logs